two mirrors are placed at an angle of 60 degree.the total no. of images of an object that can be seen are: ?

Grade:10

1 Answers

Arun
6 years ago
Dear student
 

The number of images formed in two plane mirrors inclined at an angle A to each other is given by the formula:

Number of images = 360/A - 1 = (360/60) - 1 = 6 - 1 = 5

 

Regards

Arun (askIITians forum expert)